# The rising costs of training frontier AI models

### v1

@@ Abstract
The costs of training frontier AI models have grown dramatically in recent years, but there is limited public data on the magnitude and growth of these expenses. This paper develops a detailed cost model to address this gap, estimating training costs using three approaches that account for hardware, energy, cloud rental, and staff expenses. The analysis reveals that the amortized cost to train the most compute-intensive models has grown precipitously at a rate of 2.4×2.4$\times$2.4 × per year since 2016 (95% CI: 2.0×2.0$\times$2.0 × to 3.1×3.1$\times$3.1 ×). For key frontier models, such as GPT-4 and Gemini, the most significant expenses are AI accelerator chips and staff costs, each costing tens of millions of dollars. Other notable costs include server components (15-22%), cluster-level interconnect (9-13%), and energy consumption (2-6%). If the trend of growing development costs continues, the largest training runs will cost more than a billion dollars by 2027, meaning that only the most well-funded organizations will be able to finance frontier AI models.00footnotetext: 1Epoch AI. 2Stanford University.
@@ 1 Introduction
Our methods are built upon a comprehensive database of notable machine learning models [7], and informed by interviews with industry experts. We consider three complementary approaches to measuring the cost of frontier models. The first approach estimates the hardware capital expenses (CapEx) amortized over the final training run, along with the cost of hardware energy consumption. By considering AI accelerator chips, other server hardware, networking hardware, and energy separately, this approach can provide more accurate training costs. We find that the most expensive publicly-announced training runs to date are OpenAI’s GPT-4 at \$40M and Google’s Gemini Ultra at \$30M. Among frontier models, defined as models within the top 10 most compute-intensive models when they are released, we find that training has become 2.4×2.4$\times$2.4 × more expensive per year since 2016 (95% CI: 2.0×2.0$\times$2.0 × to 3.1×3.1$\times$3.1 ×).
@@
We then compare this approach to the cloud-price approach that was first presented in the AI Index [6]. Instead of estimating hourly compute costs in detail, the cloud-price approach simply uses historical rental rates from cloud platforms. The cloud-price approach shows a similar growth rate (2.6×2.6$\times$2.6 × per year with a 95% CI of 2.1×2.1$\times$2.1 × to 3.2×3.2$\times$3.2 ×), but yields costs that are about twice as large on average. We expect the cloud-price approach to overestimate frontier model costs, since model developers usually either own or have private rental agreements for their training hardware. Using both approaches helps validate our estimate of cost growth, while also highlighting the uncertainty of individual costs.
Our third and most in-depth approach breaks down hardware, energy, and R&D staff costs over the entire development of the model (i.e. both experiments and training). We select four especially notable models for this approach—GPT-3, OPT-175B, GPT-4, and Gemini Ultra. For these models, we find that R&D staff costs including equity are between 29% and 49% of the total amortized cost. Computing hardware makes up 47–65%, while energy comprises only 2–6%. However, if we exclude equity the fraction for R&D staff drops to 22–33%, and the fractions of computing hardware costs and energy rise to 60–74% and 2–7% respectively.
@@ 2.1 Datasets and frontier model selection
Our investigation draws upon the Epoch database, which documents 796 notable models across the history of machine learning [7]. Key details captured for each model include training compute, dataset size, and parameter count. To focus on the largest-scale models, we initially filtered the database to models that had training compute estimates and that were published on or after 1 October 2015 (the start of the large-scale ML era according to [8]) and up to 31 December 2023. This resulted in 276 selected models. For these models, we recorded the training time, hardware type and quantity, and utilization rate sourced from each model’s original publication, where possible.
For our main results, we examined 45 models that were historically at the frontier of compute. Specifically, we filtered for models that were in the top 10 of training compute as of their release. Section B.1 provides further details on this selection procedure and a comparison to three alternative methods.
@@ 2.2 Amortizing the cost of hardware for training
To determine the cost of hardware for a training run, we first calculated the acquisition cost of the necessary accelerator chips (GPUs/TPUs), servers, and networking hardware. This involved looking up historical prices for GPUs, or estimating production costs for TPUs. Further details are provided in Section A.1. This hardware generally remains available for future use after a training run finishes, but its value depreciates due to both hardware failures and obsolescence caused by hardware progress. We therefore estimated depreciation and the amortized hardware acquisition cost over the training time.
We depreciated hardware value by a rate of 0.14 orders of magnitude per year, based on the median estimate of the growth rate in ML GPU price-performance in [9]. [^1] This growth rate measures improvement at 32-bit precision. One-time improvements from lower-precision and tensor number formats would make the rate faster, but this was not estimated. This also assumes that hardware improves continuously. In reality, hardware improves in increments with each new release. To estimate the depreciated value of hardware at the start of training, we used the following formula:
Initial hardware value=Hardware acquisition $\text{Initial hardware value}=\dfrac{\text{Hardware acquisition cost}}{\exp\Big{(}\big{[}\text{Training start date}-\text{Hardware availability date}\big{]}\times\text{Depreciation factor}\Big{)}}$
We neglected the impact of hardware failures on depreciation, because manufacturers such as NVIDIA have a warranty that repairs or replaces defective hardware [10]. While some cost is incurred by downtime between hardware replacements, failure rates seem low enough to make this negligible compared to depreciation due to hardware progress. We provide evidence for that in Section A.3.
One challenge is that we often know chip-hours for a given training run, but not the exact training duration, so we could not use the same depreciation model for the training run. Instead, we used a linear approximation:
[…]
where training chip-hours is the product of training time and hardware quantity. For our first two approaches we only consider the chip-hours of the final training run, but in Section 3.5 we expand this to account for all experiments towards developing the model. While the depreciation and amortization process involved several estimates and approximations, our results are robust across reasonable variations (see Section A.3 for further methodological details and Section B.3 for the sensitivity analysis).
@@ 2.3 Hardware energy consumption cost
Total energy cost of training=Energy cost rate (\$/kWh)×Hardware TDP (kW)×Average power to TDP ratio (%)×Data center PUE×Number of chip-hours ($\begin{split}\text{Total energy cost of training}=\ &\text{Energycost rate (\$/kWh)}\times\text{Hardware TDP (kW)}\ \times\\ &\text{Average power to TDP ratio (\%)}\times\text{Data center PUE}\times\text{Number of chip-hours (h)}\end{split}$
@@ 2.4 Cloud compute cost
To estimate training costs from cloud rental prices, we used the following cost model:
@@ 2.5.1 R&D staff costs
For full-time equivalent workers, we were informed by the type and number of contributors listed on the research paper. For all models except Gemini Ultra, we sampled full-time equivalent workloads from a 90% credible interval of 5% to 80% FTE for each contributor, resulting in a median of 20%. For Gemini Ultra, we used different workloads for each type of contributor listed [11, pp. 66–69].
@@ 3 Results
The amortized training costs of frontier models have increased by a factor of 2.4x per year since 2016. This is the result of the preferred amortized hardware CapEex + energy approach, shown in Figure 2. Table 1 compares this to the cloud approach, which yields a similar growth rate of 2.6×2.6$\times$2.6 × per year. The growth rate is also similar if we vary hardware depreciation or training start date within reasonable limits (see Section B.4). However, the growth rate rises to 2.9x per year if we exclude TPUs, which have more uncertain costs than publicly-sold GPUs.
@@
2.6 [2.1, 3.2] 0.41 [0.32, 0.51] 9 [7, 11] 0.68 40
Table 1: Cost growth rates based on log-linear regression, for different cost estimation approaches. All approaches select the top 10 most compute intensive models at the time of model release. $N$ refers to the number of relevant observations. Based on a two-sided $t$-test adjusted for correlation of residuals, the growth rates for amortized hardware capex + energy and cloud are not significantly different $(p=0.13)$. However, when the costs of models trained with estimated TPU production costs are excluded, the growth rate rises significantly to 2.9x per year ($p<0.01$). OOMs/year: orders of magnitude per year. Square brackets: 95% confidence interval.
@@
Estimating costs from cloud rental prices, although less representative of actual costs, has the advantage of simplicity. The cloud cost approach also helps to check the robustness of the amortized hardware CapEx + energy approach. Figure 3 shows the trend of cloud compute cost to train models among the top 10 most compute-intensive as of their release. Note that some of these estimates previously appeared in the 2024 AI Index report [6]. We find that the cost of training models based on cloud rental prices has grown by 2.6×2.6$\times$2.6 × per year since 2016, with a 95% CI of 2.1×2.1$\times$2.1 × to 3.2×3.2$\times$3.2 ×. This is consistent with the amortized hardware CapEx + energy approach, as shown in Table 1. This shows that our trend estimates are robust to two different ways of estimating prices per chip-hour.
@@
3.2 Today’s most expensive models cost tens of millions of dollars to train, implying one billion dollars by 2027
While growth rates in training cost indicate how rapidly AI investment is scaling, the actual cost of the largest frontier models is also important to grasp. Looking back to 2017 in Figure 2, AlphaGo Zero was the most expensive for its time, at \$600K. Five years later, the model with the largest compute expense is GPT-4 at \$40M, with Gemini Ultra slightly lower at \$30M. This implies that, at a growth rate of 2.4×2.4$\times$2.4 × per year, the largest training run in 2027 will cost over \$1 billion in compute expenses.
The cost of Gemini Ultra is surprisingly low given it used more training compute than GPT-4 [7]. This is partly due to Google manufacturing their own TPUs, avoiding high markups on GPUs. Another reason is depreciation: our model depreciates the value of TPU v4 hardware by more than half in the two years between the announcement of TPU v4 and the training of Gemini Ultra.
In general, the individual cost estimates are highly uncertain. A simple analysis suggests that the estimates for models trained on GPUs are only accurate within a factor of three to four (at 90% confidence; see Section A.7). The costs of models trained on TPUs are even more uncertain, being accurate within a factor of five. Despite this uncertainty in individual cost estimates, estimates of overall cost growth are more confident, as they are driven by well-characterized trends in compute growth.
@@ 3.3 Hardware acquisition costs are one to two orders of magnitude higher than amortized costs
To illustrate the difference between amortized hardware cost and acquisition cost, Figure 4 shows the acquisition costs we were able to estimate using hardware purchase prices and training hardware quantities. Since this is the up-front cost of acquiring the hardware, the costs are one to two orders of magnitude larger than amortized hardware costs. For example, we estimate that it cost \$800M to acquire the hardware used to train GPT-4, compared to \$40M for the amortized hardware CapEx + energy cost. The ratio between the two depends on when and for how long the model is trained. Based on 40 estimates of acquisition cost, we find a growth rate of 2.5×2.5$\times$2.5 × per year (95% CI: 2.0×2.0$\times$2.0 ×, 3.0×3.0$\times$3.0 ×), which is similar to the growth in amortized costs.
@@ 3.5 R&D staff are a significant fraction of costs over the whole model development process
We find that when equity is included, R&D staff costs make up between 29% and 49% of total amortized model development costs, depending on the model. Excluding equity, the fraction decreases to 19% to 33% (see Section B.5 for additional plots). Notably, this fraction does not change much from GPT-3 to GPT-4, which spans three and a half years of AI progress. The number of reported contributors increased from 25 for GPT-3 [13] to 284 for GPT-4 [14], while the amortized hardware cost over the whole model development increased from \$4M to \$90M. However, due to the limited data, we caution against extrapolating the fraction of R&D staff costs to future frontier models.
@@
On the compute side, we find that amortized hardware cost makes up 47–65% of the full model development cost, while energy comprises only 2–6%. With equity excluded from R&D costs, the fraction of hardware cost and energy cost rise to 60–74% and 2–7% respectively. Note that while energy consumption is a small fraction of total cost, this doesn’t entail that power requirements are not a challenge in frontier AI development. Regulatory and logistical hurdles to secure power supplies may cause bottlenecks in the coming years, but we leave that topic to future work.
@@ 5 Conclusion
In this paper we used three approaches to analyze the cost of training ML models at the frontier. The first two approaches—one based on hardware purchase prices and energy costs, the other based on cloud rental prices—indicate that the amortized cost of compute for these training runs has grown by around 2.4x per year (95% CI: 2.0x to 3.1x) since 2016. This shows the large role of investment in driving AI progress.
@@ Acknowledgements
We thank Bartosz Podkanowicz for assisting with data collection, Luke Frymire for assisting with ground truth cost verification, and Josh You for copyediting. We thank Igor Molybog, Yafah Edelman and Horace He for helpful conversations about the training process and requirements for creating large ML models. We thank Konstantin Pilz, Yafah Edelman, Tom Davidson, Isabel Juniewicz, Carl Shulman, Jaime Sevilla, Tamay Besiroglu, Aleksandar Kostovic, Tim Fist, Haydn Belfield, Alan Chan, David Patterson, Mauricio Baker, Erich Grunewald, and Cullen O’Keefe for feedback on drafts.

### v2

@@ Abstract
The costs of training frontier AI models have grown dramatically in recent years, but there is limited public data on the magnitude and growth of these expenses. This paper develops a detailed cost model to address this gap, estimating training costs using three approaches that account for hardware, energy, cloud rental, and staff expenses. The analysis reveals that the amortized cost to train the most compute-intensive models has grown precipitously at a rate of 2.4×2.4$\times$2.4 × per year since 2016 (90% CI: 2.0×2.0$\times$2.0 × to 2.9×2.9$\times$2.9 ×). For key frontier models, such as GPT-4 and Gemini, the most significant expenses are AI accelerator chips and staff costs, each costing tens of millions of dollars. Other notable costs include server components (15-22%), cluster-level interconnect (9-13%), and energy consumption (2-6%). If the trend of growing development costs continues, the largest training runs will cost more than a billion dollars by 2027, meaning that only the most well-funded organizations will be able to finance frontier AI models.00footnotetext: 1Epoch AI. 2Stanford University.
@@ 1 Introduction
Our methods are built upon a comprehensive database of notable machine learning models [7], and informed by interviews with industry experts. We consider three complementary approaches to measuring the cost of frontier models. The first approach estimates the hardware capital expenses (CapEx) amortized over the final training run, along with the cost of hardware energy consumption. By considering AI accelerator chips, other server hardware, networking hardware, and energy separately, this approach can provide more accurate training costs. We find that the most expensive publicly-announced training runs to date are OpenAI’s GPT-4 at \$40M and Google’s Gemini Ultra at \$30M. Among frontier models, defined as models within the top 10 most compute-intensive models when they are released, we find that training has become 2.4×2.4$\times$2.4 × more expensive per year since 2016 (90% CI: 2.0×2.0$\times$2.0 × to 2.9×2.9$\times$2.9 ×).
@@
We then compare this approach to the cloud-price approach that was first presented in the AI Index [6]. Instead of estimating hourly compute costs in detail, the cloud-price approach simply uses historical rental rates from cloud platforms. The cloud-price approach shows a similar growth rate (2.5×2.5$\times$2.5 × per year with a 90% CI of 2.1×2.1$\times$2.1 × to 3.1×3.1$\times$3.1 ×), but yields costs that are about twice as large on average. We expect the cloud-price approach to overestimate frontier model costs, since model developers usually either own or have private rental agreements for their training hardware. Using both approaches helps validate our estimate of cost growth, while also highlighting the uncertainty of individual costs.
Our third and most in-depth approach breaks down hardware, energy, and R&D staff costs over the entire development of the model (i.e. both experiments and training). We select four especially notable models for this approach—GPT-3, OPT-175B, GPT-4, and Gemini Ultra. For these models, we find that R&D staff costs including equity are between 29% and 49% of the total amortized cost. Computing hardware makes up 47–64%, while energy comprises only 2–6%. However, if we exclude equity the fraction for R&D staff drops to 19–33%, and the fractions of computing hardware costs and energy rise to 61–76% and 2–7% respectively.
@@ 2.1 Datasets and frontier model selection
Our investigation draws upon the Notable AI Models database, which documents 796 notable models across the history of machine learning [7]. Key details captured for each model include training compute, dataset size, and parameter count. To focus on the largest-scale models, we initially filtered the database to models that had training compute estimates and that were published on or after 1 October 2015 (the start of the large-scale ML era according to [8]) and up to 31 December 2023. This resulted in 276 selected models. For these models, we recorded the training time, hardware type and quantity, and utilization rate sourced from each model’s original publication, where possible.
For our main results, we examined 41 models that were historically at the frontier of compute. Specifically, we filtered for models that were in the top 10 of training compute as of their release. [^1] We excluded models that are fine-tuned versions of a separately listed model, to avoid double-counting costs. Section B.1 provides further details on this selection procedure and a comparison to three alternative methods.
@@ 2.2 Amortizing the cost of hardware for training
To estimate the cost of hardware for a training run, we first calculated the cost to acquire the necessary accelerator chips (GPUs/TPUs), servers, and networking hardware. This involved looking up historical prices for GPUs, or estimating production costs for TPUs. Further details are provided in Section A.1.
Hardware normally remains available for future use after a training run finishes, but its value depreciates over time due to hardware progress. We amortized the cost of a training run based on this depreciation. Specifically, we depreciated the value of hardware at a rate of $r=0.14$ orders of magnitude per year, based on the growth rate of ML GPU price-performance [9]. [^2] This growth rate measures improvement at 32-bit precision. One-time improvements from lower-precision and tensor number formats would make the rate faster, but this was not estimated. This also assumes that hardware improves continuously. In reality, hardware improves in increments with each new release. To get the value of the hardware at the start of training, we used the following formula:
Start value per chip=Acquisition cost per $\text{Start value per chip}=\dfrac{\text{Acquisition cost per chip}}{\exp\Big{(}\big{[}\text{Training start date}-\text{Hardware availability date}\big{]}\cdot r\ln{10}\Big{)}}$
where the difference in dates is in units of years. For example, if the training run starts one year after hardware is acquired, the start value is approximately 72% of the acquisition cost. We neglected the impact of hardware failures on depreciation, as the effect seemed small compared to hardware progress. We provide evidence for that in Section A.3.
After finding the initial value of the hardware, the amortized cost of the training run is then the portion of value that is lost during the training run. However, the training time is often difficult to determine, due to a lack of public information. More often, we were able to estimate the number of chip-hours: the product of the training time and the number of chips. So we substituted chip-hours for the training time and the number of chips, using a linear approximation. This led to our final formula for amortized training cost:
Amortized training costStart value per chip×Training chip-hours(365×24) $\text{Amortized training cost}\approx\text{Start value per chip}\times\frac{\text{Training chip-hours}}{(365\times 24)\text{ hours/year}}\times r\ln{10}$
Up until Section 3.5, our results only account for the chip-hours of the final training run. In Section 3.5, we scale up the chip-hours to account for all experiments towards developing an AI model. Although the amortized cost model involves several estimates and approximations, our results are robust to reasonable changes in the method (see Section A.3 for further methodological details and Section B.3 for the sensitivity analysis).
@@ 2.3 Hardware energy consumption cost
Total energy cost of training=Energy cost rate (\$/kWh)×Hardware TDP (kW)×Average power to TDP ratio (%)×Data center PUE×Training chip-hours ($\begin{split}\text{Total energy cost of training}=\ &\text{Energycost rate (\$/kWh)}\times\text{Hardware TDP (kW)}\ \times\\ &\text{Average power to TDP ratio (\%)}\times\text{Data center PUE}\times\text{Training chip-hours (h)}\end{split}$
@@ 2.4 Cloud compute cost
To estimate training costs from cloud rental prices, we used the following formula:
@@ 2.5.1 R&D staff costs
For full-time equivalent workers, we were informed by the type and number of contributors listed on the research paper. For all models except Gemini Ultra, we sampled full-time equivalent workloads from a 90% credible interval of 5% to 80% FTE for each contributor, resulting in a median of 20%. For Gemini Ultra, we used different workloads for each type of contributor listed [10, pp. 66–69].
@@ 3 Results
The amortized training costs of frontier models have increased by a factor of 2.4x per year since 2016. This is the result of the preferred amortized hardware CapEex + energy approach, shown in Figure 2. Table 1 compares this to the cloud approach, which yields a similar growth rate of 2.5×2.5$\times$2.5 × per year. The growth rate is also similar if we vary hardware depreciation or training start date within reasonable limits (see Section B.4). However, the growth rate rises to 2.9x per year if we exclude TPUs, which have more uncertain costs than publicly-sold GPUs.
@@
2.5 [2.1, 3.1] 0.40 [0.32, 0.48] 9 [7, 11] 0.66 36
Table 1: Cost growth rates based on log-linear regression, for different cost estimation approaches. All approaches select the top 10 most compute intensive models at the time of model release. $N$ refers to the number of relevant observations. Based on a two-sided $t$-test adjusted for correlation of residuals, the growth rates for amortized hardware capex + energy and cloud are not significantly different $(p=0.13)$. However, when the costs of models trained with estimated TPU production costs are excluded, the growth rate rises significantly to 2.9x per year ($p<0.01$). OOMs/year: orders of magnitude per year. Square brackets: 90% confidence interval.
@@
Estimating costs from cloud rental prices, although less representative of actual costs, has the advantage of simplicity. The cloud cost approach also helps to check the robustness of the amortized hardware CapEx + energy approach. Figure 3 shows the trend of cloud compute cost to train models among the top 10 most compute-intensive as of their release. Note that some of these estimates previously appeared in the 2024 AI Index report [6]. We find that the cost of training models based on cloud rental prices has grown by 2.5×2.5$\times$2.5 × per year since 2016, with a 90% CI of 2.1×2.1$\times$2.1 × to 3.1×3.1$\times$3.1 ×. This is consistent with the amortized hardware CapEx + energy approach, as shown in Table 1. This shows that our trend estimates are robust to two different ways of estimating prices per chip-hour.
@@
3.2 The trend suggests that the most expensive publicly announced model will cost one billion dollars to train by the start of 2027
The growth rate in training cost indicates how rapidly AI investment is scaling. We can use this growth rate to extrapolate the cost of the largest training run. Currently, GPT-4 has the largest amortized hardware and energy cost, at \$40M. GPT-4 was published in March of 2023 [12]. This implies that, at a growth rate of 2.4×2.4$\times$2.4 × per year, the most expensive publicly announced model by the start of 2027 will cost about \$1 billion.
Whether this cost is justified hinges on how profitable the resulting AI model is—but parts of the AI industry believe it is worthwhile. The CEO of the AI lab Anthropic has claimed that close to a billion dollars will already be spent on a single training run in 2024 (implying an amortized cost), which is even sooner than the historical trend suggests [5].
@@ 3.3 Hardware acquisition costs are one to two orders of magnitude higher than amortized costs
To illustrate the difference between amortized hardware cost and acquisition cost, Figure 4 shows the acquisition costs we were able to estimate using hardware purchase prices and training hardware quantities. Since this is the up-front cost of acquiring the hardware, the costs are one to two orders of magnitude larger than amortized hardware costs. For example, we estimate that it cost \$800M to acquire the hardware used to train GPT-4, compared to \$40M for the amortized hardware CapEx + energy cost. The ratio between the two depends on when and for how long the model is trained.
Based on 40 estimates of acquisition cost, we find a growth rate of 2.5×2.5$\times$2.5 × per year (90% CI: 2.1×2.1$\times$2.1 ×, 3.0×3.0$\times$3.0 ×). This is slightly faster than the rate of 2×2$\times$2 × per year suggested by amortized costs (2.4×2.4$\times$2.4 × per year) divided by training times (1.2×1.2$\times$1.2 × per year) [13]. The discrepancy is due to different AI models appearing in each analysis, highlighting a source of sensitivity in our results.
@@ 3.5 R&D staff are a significant fraction of costs over the whole model development process
We find that when equity is included, R&D staff costs make up between 29% and 49% of total amortized model development costs, depending on the model. Excluding equity, the fraction decreases to 21% to 33% (see Section B.5 for additional plots). Notably, this fraction does not change much from GPT-3 to GPT-4, which spans three and a half years of AI progress. The number of reported contributors increased from 25 for GPT-3 [14] to 284 for GPT-4 [12], while the amortized hardware cost over the whole model development increased from \$4M to \$90M. However, due to the limited data, we caution against extrapolating the fraction of R&D staff costs to future frontier models.
@@
On the compute side, we find that amortized hardware cost makes up 47–64% of the full model development cost, while energy comprises only 2–6%. With equity excluded from R&D costs, the fraction of hardware cost and energy cost rise to 61–76% and 2–7% respectively. Note that while energy consumption is a small fraction of total cost, this doesn’t entail that power requirements are not a challenge in frontier AI development. Regulatory and logistical hurdles to secure power supplies may cause bottlenecks in the coming years, but we leave that topic to future work.
@@ 5 Conclusion
In this paper we used three approaches to analyze the cost of training ML models at the frontier. The first two approaches—one based on hardware purchase prices and energy costs, the other based on cloud rental prices—indicate that the amortized cost of compute for these training runs has grown by around 2.4x per year (90% CI: 2.0x to 2.9x) since 2016. This shows the large role of investment in driving AI progress.
@@ Acknowledgements
We thank Bartosz Podkanowicz for assisting with data collection, Luke Frymire for assisting with ground truth cost verification, and Josh You for copyediting. We thank Igor Molybog, Yafah Edelman and Horace He for helpful conversations about the training process and requirements for creating large ML models. We thank Konstantin Pilz, Yafah Edelman, Tom Davidson, Isabel Juniewicz, Carl Shulman, Jaime Sevilla, Aleksandar Kostovic, Tim Fist, Haydn Belfield, Alan Chan, David Patterson, Mauricio Baker, Erich Grunewald, and Cullen O’Keefe for feedback on drafts.